6. The Evanec Company's next expected dividend, D1, is $2.62; its growth rate is 7%; and...

6.

The Evanec Company's next expected dividend, D1, is $2.62; its growth rate is 7%; and its common stock now sells for $37.00. New stock (external equity) can be sold to net $33.30 per share.

  1. What is Evanec's cost of retained earnings, rs? Do not round intermediate calculations. Round your answer to two decimal places.

    rs =   %

  2. What is Evanec's percentage flotation cost, F? Round your answer to two decimal places.

    F =   %

  3. What is Evanec's cost of new common stock, re? Do not round intermediate calculations. Round your answer to two decimal places.

    re =   %
Answer #1

a)

cost of retained earnings

= D1/price + growth rate

= 2.62/37 + 0.07

= 14.08

b)

Flotation cost = (37 - 33.3)/37

= 10%

c)

cost of common stock

= D1/price*(1-flotation cost) + growth rate

= 2.62/33.3 + 0.07

= 14.87%